Two charges +9e and +e are kept 16cm apart. Where should be the third charge q be placed between them so that system remains equillibrium

Grade:12

3 Answers

Arun
5 years ago
Given charges: +e and +9e
Let the third chare q be placed at a distance x from e and hence its distance from 9e is 16-x cm
For the system to remain in equilibrium,
F1 = F2
Therefore,
F1 = K(e q)/x2
F2 = K(9e)/(16-x)2
Equating F1 and F2 , we get,
9/x2 = 1/(16-x)2
Taking square root on both sides:
3/x = 1/(16-x)
Solving for x, we get,
x = 12cm
hence the charge q should be placed at 12 cm from e.
Naina gill
5 years ago
Two point charges +9e and +1e are kept at a distance of 16cm apart from each other at what point between these charges should a third charge q' be placed.so that it remains in equilibrium?
 
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

+e and +9e
Let the third chare q be placed at a distance x from e and hence its distance from 9e is 16-x cm
For the system to remain in equilibrium,
F1 = F2
Therefore,
F1 = K(e q)/x2
F2 = K(9e)/(16-x)2
Equating F1 and F2 , we get,
9/x2= 1/(16-x)2
Taking square root on both sides:
3/x = 1/(16-x)
Solving for x, we get,
x = 12cm
hence the charge q should be placed at 12 cm from e.

Thanks and Regards